# A Multi-agent LLM-based JUnit Test Generation with Strong Oracles

### v2

@@ 3.1 Step I: Initialization
The Initializer is a basic LLM agent responsible for producing an initial test file based on the given source code. As illustrated in the first profile in Figure 3, the Initializer is guided by two types of prompts: a system prompt, which defines the model’s role (e.g., as a software testing expert), and a user prompt, which provides specific inputs such as the source code, last failed test file, error messages, Java and JUnit 5 conventions and output format. The Initializer responds with a JSON object containing a single field, test_file, which holds the complete generated test code. Given the source file provided in Section 2, we present the generated initialize test file under Profile 1 of Figure 3, including one test case testOneElement.
@@
The generated test file is then passed to the Validation component, which attempts to compile and execute it. At this stage, often contains syntactic errors and achieves low code coverage. In the example of showed in Figure 2, the input list variable is incorrectly instantiated as List<Object> input = [1] following Python instead of Java conventions. If syntax-related errors are detected, the error messages are fed back to the Initializer for refinement. This process is repeated until a syntactically correct and error-free test file is produced or a maximum number of attempts (max_attempts) is reached. An example of is shown in Figure 2, where the input list variable is correctly instantiated as List<Object> input = ArrayList.asList(1) following Java conventions. While is syntactically valid and adheres to language and testing framework conventions, it typically still has low coverage, relying on subsequent steps for coverage improvement.
3.2 Step II: Test Prefix Generation
@@
The Planner is a basic LLM agent responsible for analyzing coverage gaps and proposing new test cases accordingly. As shown in the second profile in Figure 3, the system prompt defines its role as a test plan designer tasked with improving code coverage. The user prompt requires inputs including the source code, uncovered lines from the Jacoco coverage report, and the current test file. Based on this information, the Planner outputs a list of structured test plans in JSON format under the field test_cases_to_add, with each test case described by a name, a brief natural language description, input, and expected output. A snippet of a generated plan based on the running example is provided in Profile 2, Figure 3. This plan proposes two test cases to enhance code coverage: Test empty list and Test list with multiple of 4 not 3 index.
@@
The Tester is a basic LLM agent that generates executable test code based on the plans proposed by the Planner. As shown in the third profile in Figure 3, the system prompt defines the Tester’s role as a Java unit test generator following JUnit 5 and the conventions of the given initial test file. The user prompt includes the source code, the test plan from the Planner, and feedback from the Inspector (if any). The output is a JSON object containing a list of generated_test_cases, with each entry specifying the behavior—a natural language explanation of what the test is verifying, test_name, test_code, and any new_import_statements required to execute this test case. A snippet of the generated tests based on the running example is provided under Profile 3 in Figure 3. The Tester generates the test case testMultipleOf4Not3, with a test prefix of ArrayList.asList(1,2,3,4,5) and a tentative oracle assertEquals(27, SumSquares1.sumSquares(input)). Notice the tentative oracle is incorrect: it is calculated as , whereas the correct oracle should be . The fifth element was not cubed as specified in the description due to a bug in the source code, specifically the conditional statement if (i%4==0&&i%3==0) (Figure 1). Additionally, the Tester fails to identify the necessary import statements for ArrayList.
@@
The Inspector is a basic LLM agent that evaluates the compiled and executed test cases to detect and explain any errors. Its role, as defined in the fourth profile in Figure 3, is to inspect the generated test code and offer feedback for improvements. The user prompt includes the source code and error messages from the validation process. The output consists of structured feedback entries, each containing the failed_test_code, the error_message, the error_type, and a potential_fix. This feedback is then sent back to the Tester for refinement. A snippet of the generated feedback based on the running example is presented under Profile 4 of Figure 3. In this case, the Inspector correctly identifies the missing import statements for ArrayList in the test file . However, it does not flag the incorrect oracle, as the analysis is performed against the faulty source code, which exhibits the same buggy behavior as the test output.
@@
3.3 Step III: Oracle Fixing
@@
The Requirement Engineer is a basic LLM agent responsible for extracting both natural language requirements and formal specifications from the provided description of the function. As illustrated in the fifth profile in Figure 3, the system prompt defines its role as a software engineer specializing in requirement analysis. The user prompt includes the textual description of the code, and the output consists of a set of human-readable requirements (e.g., input type and expected behavior). When possible, it also produces a formal specification expressed in predicate logic. If a formalization is not applicable or too ambiguous for the task, the specification is left empty, and only the natural language requirements are used for downstream oracle correction. The extracted requirements and specifications for the running example are shown under Profile 5 in Figure 3.
@@
Each pipeline comprises a reasoning LLM agent Panelist and a basic LLM agent Interpreter. Each Panelist functions as an independent oracle corrector. As shown in the sixth profile in Figure 3, the system prompt defines its role as a testing expert tasked with identifying and fixing incorrect test oracles generated by automated tools. The user prompt requires the natural language description, the extracted requirements and specifications from the Requirement Engineer, and the set of test cases generated in Step II. Each agent analyzes the given test case and determines whether the existing oracle is correct. If not, it provides a corrected oracle based on the intended behavior described in the requirements and specification. To mitigate the “overthinking phenomenon” (see Section 1) of reasoning LLMs, each panelist is paired with a basic LLM agent Interpreter. Each Interpreter is responsible for extracting relevant insights from Panelist’s verbose thoughts and produce structured evaluations of the tentative oracles. As shown in the seventh profile of Figure 3, the system prompt defines its role as an assistant working with an excellent but self-doubting tester. The user prompt requires the Panelist’s thoughts and the test code under examination. Example output of both Panelist and Interpreter based on the running example are shown under Profile 6 and Profile 7 of Figure 3, respectively. In this case, the Panelist generates a lengthy, self-questioning reasoning process — using phrases like Wait and However — while the Interpreter successfully extracts concise and structured oracle evaluations.
@@
Multiple dual-LLM pipelines operate in parallel, each making independent judgments of the tentative oracle. The Curator aggregates these judgments and determines the final oracle. As shown in the eighth profile of Figure 3, the system prompt defines its role as a leader of a software engineering team. While teams members are discussing the correctness of the tentative oracle, the Curator is responsible for summarizing and making a final judgment. Its user prompt requires individual judgment from each pipeline, and the test code under examination. An example output of Curator based on the running example is shown below Profile 8 of Figure 3. In this case, the Curator successfully identifies the tentative oracle as incorrect and revises it into the correct oracle assertEquals(147, SumSquares1.sumSauares(input));.
@@ 4 Experimental Setup
What is the individual effectiveness contribution of key components in CANDOR, including the Planner, Requirement Engineer and the panel discussion?
@@
RQ1 evaluates the overall effectiveness of the generated test cases by comparing code coverage and mutation scores with state-of-the-art baselines. We conduct experiments on the benchmark HumanEvalJava dataset Athiwaratkun et al. (2022). Additionally, we evaluate CANDOR on a LeetCode-derived dataset (LeetCodeJava) that addresses problems of varying difficulty to assess its robustness across different levels of code complexity. For RQ2, we focus on the accuracy of oracle generation. We evaluate oracle correctness under two conditions: (1) when the source code is correct and conforms to the intended requirements, and (2) when the code is faulty. To simulate faulty scenarios, we adopt a commonly-used tool (PiTest 202 (Accessed: 2025b)) to inject mutations into the source code from both HumanEvalJava and LeetCodeJava. This evaluation investigates CANDOR’s robustness and its practical utility in real-world testing scenarios. With RQ3, we assess the individual contributions of key components within CANDOR through a series of ablation studies. These studies focus on the Planner, Requirement Engineer, and the panel discussion. We do not perform ablation on other agents, as they are essential for generating valid test cases; removing them would render CANDOR non-functional.
4.1 Datasets
In this work, we focus on unit test generation for Java. We evaluate CANDOR on two datasets: HumanEvalJava Athiwaratkun et al. (2022) and LeetCodeJava 202 (Accessed: 2025a). We exclude other datasets used in the literature such as Defects4J Just et al. (2014), SF110 Fraser and Arcuri (2011a), and OracleEval25 Hossain and Dwyer (2024), as they are project-level datasets that involve complex dependencies and user-defined classes. Incorporating such datasets requires additional handling of project-specific configurations, which is beyond the scope of this work. Moreover, these datasets do not consistently provide natural language descriptions of SUT—an essential component for our method, which relies on such descriptions to enable requirement extraction and oracle generation. We leave the integration and evaluation on project-level datasets for future work.
@@ 4.1 Datasets
It consists of 160 Java programs implemented by Siddiq et al. to solve problems introduced in Athiwaratkun et al. (2022). Each program defines a Java class under test that includes a single method. The length of these programs ranges from 20 to 222 lines, with an average of 41 lines of code.
@@
It includes solutions to programming problems, categorized by difficulty levels: “easy”, “medium”, and “hard”. More difficult problems generally require more complex code. Due to the constraints of time and available computational resources, we randomly sampled 50 programs from “medium”, and “hard” difficulty category, resulting in a total of 100 programs. These programs have an average of 33 lines of code, with a minimum of 7 and a maximum of 126. The average lines of code for “medium”, and “hard” categories are 31, and 39, respectively.
4.2 Baselines
We compare CANDOR against two representative baselines: EvoSuite Fraser and Arcuri (2011a) and LLM-Empirical Siddiq et al. (2024).
@@ 4.2 Baselines
This is a widely adopted automated test generation tool for Java programs that uses search-based techniques, such as genetic algorithms, to systematically generate unit tests. It analyzes the program under test to produce test cases aiming to maximize code coverage metrics like line coverage. In addition to generating test prefixes, EvoSuite automatically produces regression oracles—assertions that capture the current program behavior by checking expected outputs, exceptions, and state changes. These generated tests are structured as JUnit test cases, making EvoSuite a popular baseline for research in automated testing and regression testing Fraser and Arcuri (2011a); Hossain and Dwyer (2024); Tufano et al. (2020).
@@
The LLM-Empirical method, proposed by Siddiq et al. Siddiq et al. (2024), leverages LLMs such as Codex, GPT-3.5-Turbo, and StarCoder to generate JUnit tests for Java programs using prompt engineering without fine-tuning. Unlike EvoSuite, which generates regression oracles based on the current implementation behavior, LLM-Empirical generates test oracles based on natural language descriptions of the SUT. It represents the SOTA in prompt-based, LLM-driven oracle generation for Java unit tests. Among the models used in LLM-Empirical, Codex (specifically, code-davinci-002) and GPT-3.5-Turbo are the most capable, but both are closed-source. Moreover, code-davinci-002 is no longer publicly available, and GPT-3.5-Turbo requires paid API access. As a result, for HumanEvalJava, we report the results from their original paper, and for LeetCodeJava, we report the results obtained using GPT-3.5-Turbo. Other advanced LLMs cannot be directly applied to LLM-Empirical because their implementation—particularly the component that fixes test cases—is specifically tailored to the behavior and outputs of the LLM they use.
We are also aware of TOGLL Hossain and Dwyer (2024), a recent oracle generation method proposed by Hossain et al. However, we exclude it from our evaluation for two main reasons: (1) it relies on EvoSuite-generated test prefixes and LLM fine-tuning, which differs from the objectives of our end-to-end, prompt-only approach; (2) it was published recently, and due to time and resource constraints, we were unable to implement it on our datasets. We leave evaluation with TOGLL for future work.
4.3 Evaluation Metrics and Statistical Testing
We evaluate the effectiveness of CANDOR for test generation in terms of code coverage, mutation score, and oracle correctness. To reduce the influence of randomness, we repeat all the experiments three times and calculate the average of each metric. For statistical testing, we compare these metrics between CANDOR and each baseline across programs using a paired non-parametric test to assess significance.

### v6

@@ 3.1. Step I: Initialization
The Initializer is a basic LLM agent responsible for producing an initial test file based on the given source code. As illustrated in profile ① in Figure 3, the Initializer is guided by two types of prompts: a system prompt, which defines the model’s role (e.g., as a software testing expert), and a user prompt, which provides specific inputs such as the source code, last failed test file, error messages, Java and JUnit 5 conventions and output format. The Initializer responds with a JSON object containing a single field, test_file, which holds the complete generated test code. Given the source file provided in Section 2, we present the generated initialize test file under profile ① of Figure 3, including one test case testOneElement.
@@
The generated test file is then passed to the Validation component, which attempts to compile and execute it. At this stage, often contains syntactic errors and achieves low code coverage. In the example of shown in Figure 2, the input list variable is incorrectly instantiated as List<Object> input = [1] following Python instead of Java conventions. If syntax-related errors are detected, the error messages are fed back to the Initializer for refinement. This process is repeated until a syntactically correct and error-free test file is produced or a maximum number of attempts (max_attempts) is reached. An example of is shown in Figure 2, where the input list variable is correctly instantiated as List<Object> input = ArrayList.asList(1) following Java conventions. While is syntactically valid and adheres to language and testing framework conventions, it typically still has low coverage and relies on subsequent steps to improve it.
Note that the Initialization Step is optional in CANDOR. If an initial test case is already provided, CANDOR skips Step I and proceeds directly to generating test prefixes (Step II) and oracles (Step III).
3.2. Step II: Test Prefix Generation
@@
The Planner is a basic LLM agent responsible for analyzing coverage gaps and proposing new test cases accordingly. As shown in profile ② in Figure 3, the system prompt defines its role as a test plan designer tasked with improving code coverage. The user prompt requires inputs, including the source code, uncovered lines from the Jacoco coverage report, and the current test file. Based on this information, the Planner outputs a list of structured test plans in JSON format under the field test_cases_to_add, with each test case described by a name, a brief natural language description, input, and expected output. A snippet of a generated plan based on the running example is provided in profile ②, Figure 3. This plan proposes two test cases to enhance code coverage: Test empty list and Test list with multiple of 4 not 3 index.
@@
The Tester is a basic LLM agent that generates executable test code based on the plans proposed by the Planner. As shown in profile ③ in Figure 3, the system prompt defines the Tester’s role as a Java unit test generator following JUnit 5 and the conventions of the given initial test file. The user prompt includes the source code, the test plan from the Planner, and feedback from the Inspector (if any). The output is a JSON object containing a list of generated_test_cases, with each entry specifying the behavior—a natural language explanation of what the test is verifying, test_name, test_code, and any new_import_statements required to execute this test case. A snippet of the generated tests based on the running example is provided under profile ③ in Figure 3. The Tester generates the test case testMultipleOf4Not3, with a test prefix of ArrayList.asList(1,2,3,4,5) and a tentative oracle assertEquals(27, SumSquares1.sumSquares(input)). Notice the tentative oracle is incorrect: it is calculated as , whereas the correct oracle should be . The fifth element was not cubed as specified in the description due to a bug in the source code, specifically the conditional statement if (i%4==0&&i%3==0) (Figure 1). Additionally, the Tester fails to identify the necessary import statements for ArrayList.
@@
The Inspector is a basic LLM agent that evaluates the compiled and executed test cases to detect and explain any errors. Its role, as shown in profile ④ in Figure 3, is to inspect the last generated test code and its execution result and provide feedback for further improvement. The user prompt includes the source code and error messages from the validation process. The output consists of structured feedback entries, each containing the failed_test_code, the error_message, the error_type, and a potential_fix. This feedback is then sent back to the Tester for refinement. A snippet of the generated feedback based on the running example is presented under profile ④ of Figure 3. In this case, the Inspector correctly identifies the missing import statements for ArrayList in the test file . However, it does not flag the incorrect oracle, as the analysis is performed against the faulty source code, which exhibits the same buggy behavior as the test output.
@@
3.3. Step III: Oracle Fixing
@@
The Requirement Engineer is a basic LLM agent responsible for extracting both natural language requirements and formal specifications from the provided description of the function. As illustrated in profile ⑤ in Figure 3, the system prompt defines its role as a software engineer specializing in requirement analysis. The user prompt includes the textual description of the code, and the output consists of a set of human-readable requirements (e.g., input type and expected behavior). When possible, it also produces a formal specification expressed in predicate logic. The grammar follows standard predicate logic and, if a formalization is not applicable or too ambiguous for the task, the specification is left empty, and only the natural language requirements are used for downstream oracle correction. The Requirement Engineer Agent attempts to extract the specifications for every method under test and decides by itself on whether or not to provide them. The extracted requirements and specifications for the running example are shown under profile ⑤ in Figure 3.
@@
Each pipeline comprises a reasoning LLM agent Panelist and a basic LLM agent Interpreter. Each Panelist functions as an independent oracle corrector. As shown in profile ⑥ in Figure 3, the system prompt defines its role as a testing expert tasked with identifying and fixing incorrect test oracles generated by automated tools. The user prompt requires the natural language description, the extracted requirements and specifications from the Requirement Engineer, and the set of test cases generated in Step II. Each agent analyzes the given test case and determines whether the existing oracle is correct. If not, it provides a corrected oracle based on the intended behavior described in the requirements and specification. To mitigate the “overthinking phenomenon” (see Section 1) of reasoning LLMs, each Panelist is paired with a basic LLM agent Interpreter. Each Interpreter is responsible for extracting relevant insights from Panelist’s verbose thoughts and producing structured evaluations of the tentative oracles. As shown in profile ⑦ of Figure 3, the system prompt defines its role as an assistant working with an excellent but self-doubting tester. The user prompt requires the Panelist’s thoughts and the test code under examination. Example outputs of both Panelist and Interpreter, based on the running example, are shown under profiles ⑥ and ⑦ of Figure 3, respectively. In this case, the Panelist generates a lengthy, self-questioning reasoning process — using phrases like Wait and However — while the Interpreter successfully extracts concise and structured oracle evaluations.
@@
Multiple dual-LLM pipelines operate in parallel, each making independent judgments of the tentative oracle. The Curator aggregates these judgments and determines the final oracle. As shown in profile ⑧ of Figure 3, the system prompt defines its role as a leader of a software engineering team. While team members discuss the correctness of the tentative oracle, the Curator is responsible for summarizing and issuing a final judgment. Its user prompt requires individual judgment from each pipeline and the test code under examination. An example output of Curator based on the running example is shown below profile ⑧ of Figure 3. In this case, the Curator successfully identifies the tentative oracle as incorrect and revises it into the correct oracle assertEquals(147, SumSquares1.sumSquares(input));.
@@ 4. Experimental Setup
What is the individual effectiveness contribution of key components in CANDOR, including the Planner, the Requirement Engineer, and the panel discussion?
@@
RQ1 evaluates the overall effectiveness of the generated test cases by comparing code coverage and mutation scores with state-of-the-art baselines. We conduct experiments on the benchmark HumanEvalJava dataset (Athiwaratkun2023Multi-lingualModels), which is extensively studied in the literature (jiang2023impact; Siddiq2024UsingStudy; yuan; kang2025explainable; li2024hybrid; endres2024can; tian2025fixing). Additionally, we constructed a new dataset LeetCodeJava from LeetCode Platform (2025LeetCodeDataset) to assess CANDOR’s robustness across different levels of code complexity. For RQ2, we focus on the accuracy of oracle generation. We evaluate oracle correctness under two conditions: (1) when the source code is correct and conforms to the intended requirements, and (2) when the code is faulty. To simulate faulty scenarios, we use a commonly used tool (PiTest (2025PitestWebsite)) to inject mutations into the source code of both HumanEvalJava and LeetCodeJava to investigate CANDOR’s robustness and its practical utility in real-world testing scenarios. Specifically, we generated three mutants for each SUT. With RQ3, we assess the individual contributions of key components within CANDOR through a series of ablation studies. Each ablation study compares CANDOR with a variant that omits one key component, including the Planner (w/o Planner), Requirement Engineer (w/o Req.), and the panel discussion (w/o Panel and w/ Voting). In particular, “w/o Panel” retains only one Panelist and removes the Curator for oracle generation. “w/ Voting” replaces the panel discussion with a simple majority-voting mechanism, removing the Curator from CANDOR. Instead of the Curator reasoning over and summarizing the arguments of multiple Panelists, the final oracle is directly selected based on majority agreement. We do not perform ablations on other agents, as they are essential for generating valid test cases; removing them would render CANDOR non-functional. To mitigate the influence of randomness, all the experiments are repeated three times.
4.1. Datasets
In this work, we focus on unit test generation for Java. But at this stage, we consider only Java methods that do not depend on user-defined classes or external libraries, thus excluding datasets such as Defects4J (defects4j) and SF110 (evosuite2014). Addressing such dependencies requires additional techniques, such as retrieval-augmented generation (RAG) and mocking, which we plan to address in the future. As the first multi-agent-based framework for Java projects, CANDOR lays the foundation of testing Java methods using only off-the-shelf LLMs, a significant challenge in itself.
Specifically, we evaluate CANDOR on two representative datasets: HumanEvalJava (Athiwaratkun2023Multi-lingualModels) and LeetCodeJava. HumanEvalJava is a representative dataset for Java unit test generation, actively studied by recent researches (jiang2023impact; Siddiq2024UsingStudy; chatassert; Chen2021EvaluatingCode; kang2025explainable; codegeneratione; li2024hybrid; tian2025fixing; endres2024can). Further, we constructed a more challenging dataset, LeetCodeJava, to evaluate CANDOR under scenarios with different complexity. To demonstrate the representativeness of our selected datasets, we report key statistics in the remainder of this section, including the number of programs under test, the average line of code, and cyclomatic complexity (CC). We remark that CC is a well-established measure of control-flow complexity, quantifying the number of linearly independent execution paths in a method (cc1; cc2; cc3; bajeh2020object). Higher CC implies more branches to explore, thus making both test prefix and oracle generation more challenging. According to a large-scale empirical study of more than Java classes from open GitHub repositories (ccpaper), the average CC of Java classes is .
@@ 4.1. Datasets
It consists of 160 Java programs implemented by Siddiq et al. to solve problems introduced in (Athiwaratkun2023Multi-lingualModels). Each program defines a Java class under test containing a single method. The length of these programs ranges from 20 to 222 lines, with an average of 41 lines of code. The average CC of HumanEvalJava methods is 4.90, which is close to the average CC for Java methods (5.46) reported in (cc). This indicates that HumanEvalJava exhibits comparable control-flow complexity and is representative of typical Java methods.
@@
The LeetCode Platform lists over programming problems, categorized by difficulty levels: “easy”, “medium”, and “hard” (2025LeetCodeDataset). More difficult problems generally require more complex code. Due to time and computational resource constraints, we randomly sampled 50 programming problems from the ”medium” and ”hard” difficulty categories, for a total of 100. Then, we construct the LeetCodeJava dataset by collecting their corresponding solutions (leetcoderepo). These solutions are considered correct implementations because they are actively maintained and validated by the community. In addition to these practical considerations, we specifically chose programs of greater difficulty to ensure the dataset included more complex structural characteristics, thereby providing a more rigorous evaluation of our approach. These methods have an average of 33 lines of code, ranging from 7 to 126. The average number of lines of code for the “Medium” and “Hard” categories is 31 and 39, respectively. The average CC of methods in LeetCodeJava is generally higher than that of HumanEvalJava (4.90), with CC values for LeetCode-Medium and LeetCode-Hard reaching 6.30 and 9.46, respectively. These statistics highlight that LeetCodeJava-Medium includes Java classes with CC levels comparable to those observed in real-world code () (ccpaper), making it representative, while LeetCodeJava-Hard also contains more challenging cases with higher CC values.
4.2. Baselines
We compare CANDOR against four representative baselines: EvoSuite (Fraser2011EvoSuite:Software), LLM-Empirical (Siddiq2024UsingStudy), TOGLL (Hossain2024TOGLL:LLMs) and EVO-CANDOR. To the best of our knowledge, CANDOR is the first end-to-end, multi-agent-based approach for Java test generation that jointly addresses both test prefix and specification-based oracle generation. Among the baselines, LLM-Empirical is the most closely related, as it also performs end-to-end test generation but relies on a single off-the-shelf LLM. In contrast, EvoSuite and TOGLL represent state-of-the-art techniques for test prefix generation and specification-based oracle generation, respectively.
@@ 4.2. Baselines
This widely adopted automated test-generation tool for Java programs uses search-based techniques, such as genetic algorithms, to systematically generate unit tests. It analyzes the program under test to generate test cases that maximize code coverage metrics, such as line and branch coverage. In addition to generating test prefixes, EvoSuite automatically produces regression oracles—assertions that capture the current program behavior by checking expected outputs, exceptions, and state changes. These generated tests are structured as JUnit test cases, making EvoSuite a popular baseline for research in automated testing and regression testing (Fraser2011EvoSuite:Software; Hossain2024TOGLL:LLMs; Tufano2021UnitContext). As recently reported (tang2024chatgpt), EvoSuite remains the SOTA approach in generating test prefixes in terms of code coverage, even when compared to the latest LLM-based approaches. EvoSuite thus serves as a baseline for test prefix generation in this work.
@@
The LLM-Empirical method, proposed by Siddiq et al. (Siddiq2024UsingStudy), leverages LLMs such as Codex, GPT-3.5-Turbo, and StarCoder to generate JUnit tests for Java programs using prompt engineering without fine-tuning. Unlike EvoSuite, which generates regression oracles based on the current implementation behavior, LLM-Empirical generates test oracles based on natural language descriptions of the SUT. It represents the SOTA in prompt-based, LLM-driven oracle generation for Java unit tests. Among the models used in LLM-Empirical, Codex (specifically, code-davinci-002) and GPT-3.5-Turbo are the most capable. Moreover, code-davinci-002 is no longer publicly available. As a result, we implement LLM-Empirical using GPT-3.5-Turbo. Other advanced LLMs cannot be directly applied to LLM-Empirical because their implementation—particularly the component that fixes test cases—is specifically tailored to the behavior and outputs of the LLM they use. EvoSuite serves as a baseline for both test prefix and oracle generation in this work.
TOGLL
Recently, Hossain2024TOGLL:LLMs prensented TOGLL, fine-tuning LLMs on the SF110 dataset to generate specification-based oracles. Experimental results demonstrate that TOGLL outperforms prior oracle generators by large margins in terms of oracle correctness, bug-detection effectiveness, and mutation score, setting a new SOTA in JUnit oracle generation. Additional experiments show that TOGLL achieves optimal performance by fine-tuning a CodeParrot LLM with a prompt template that includes both program descriptions and method signatures. However, since method signatures are neither commonly available in practice nor in our experimental datasets, we do not include them in the prompt. Specifically, the prompt templates include a program description, program source code, and a test prefix generated by EvoSuite. We note that comparing TOGLL and CANDOR is biased against the latter, as TOGGL is fine-tuned on external datasets, whereas CANDOR uses off-the-shelf, general-purpose LLMs. Nevertheless, we list TOGLL as a baseline in this work, given its SOTA performance in the oracle generation domain, while EvoSuite serves as a baseline for test oracle generation.
EVO-CANDOR
TOGLL generates oracles based on EvoSuite test prefixes, while CANDOR conducts end-to-end generation, with both test prefixes and oracles generated by LLMs. Oracles for different test prefixes are not directly comparable. To ensure a fair comparison of TOGLL and CANDOR in terms of oracle quality (RQ2), we consider a variant of CANDOR, EVO-CANDOR, which generates test oracles based on the identical EvoSuite test prefixes used by TOGLL.
4.3. Evaluation Metrics and Statistical Testing
We evaluate the effectiveness of CANDOR for test generation in terms of code coverage, mutation score, and oracle correctness. The coverage and mutation scores are calculated only for test prefixes with regression oracles to ensure a fair comparison with EvoSuite. To reduce the influence of randomness, we repeat all the experiments three times and calculate the average of each metric. We also evaluate the significance of differences between CANDOR and baselines by conducting statistical tests. Specifically, we compare these metrics between CANDOR and each baseline across programs using a paired non-parametric statistical test (i.e., Wilcoxon Signed Rank test).